# Dexterous Manipulation through Imitation Learning: A Survey

### v3

@@ VI Challenges and Future Directions in Imitation Learning-Based Dexterous Manipulation
Imitation learning-based dexterous manipulation poses unique challenges due to the inherent complexities of both IL and dexterous control. Despite significant advancements over the past decade, several challenges hinder its human-level dexterity and real-world applicability. This section discusses these challenges from multiple perspectives and explores future research directions.
@@ VI-A1 Heterogeneous Data Fusion
Dexterous manipulation relies on multi-modal sensory inputs (e.g. visual, tactile, proprioceptive, and force), each with varying sampling rates, noise characteristics, and spatial-temporal resolutions, making data integration and synchronization challenging. Moreover, differences in embodiments and gripper designs introduce additional complexities. For example, demonstrations collected with one robotic hand may not directly generalize well to another due to variations in kinematics, actuation mechanisms, and sensor placements. Addressing these challenges requires (1) multi-modal alignment techniques to improve sensor fusion and (2) cross-embodiment learning frameworks for better transferability across robotic platforms and embodiments.
@@ VI-C Generalization to Novel Setups
Generalizing imitation learning-based dexterous manipulation policies is challenging due to task and environment variability, adaptive learning limitations, sim-to-real transfer issues, and cross-embodiment adaptability:
@@ VI-C2 Adaptive and Continual Learning Frameworks
VI-C3 Sim-to-Real Transfer
While simulation provides a scalable and controlled environment for training dexterous manipulation policies, transferring these learned behaviors to real-world settings is challenging. Differences in contact dynamics, sensor noise, actuation delays, and material properties create a sim-to-real gap, causing trained models to perform inconsistently when deployed on real robots. Potential future research directions are improving the realism of physics simulations through differentiable physics engines, adaptive parameter tuning, and self-supervised real-to-sim refinement to better approximate real-world interactions. Additionally, leveraging hybrid learning approaches, where policies are pre-trained in simulation and fine-tuned with real-world corrections, can enhance transferability. Uncertainty estimation techniques can also be integrated to help models recognize and adapt to distribution shifts when deployed in unstructured real-world environments.
VI-C4 Cross-Embodiment Adaptability
@@
Variability in robot embodiments, gripper designs, sensor configurations, and actuation dynamics pose significant challenges for generalization. A policy trained on one robotic hand may struggle to transfer to another due to differences in degrees of freedom, joint limits, contact dynamics, and control strategies. Even within the same robotic platform, inconsistencies arise from sensor noise, latency, and mechanical tolerances. To address this, morphology-agnostic policy learning can be explored, where models are trained across diverse robotic embodiments to develop transferable representations. Graph-based and latent-space embeddings of robot kinematics could help policies reason about different embodiments more effectively. Additionally, modular policy architectures, where separate components (e.g., perception, control, and adaptation modules) are fine-tuned independently, may enhance transferability. Another promising direction is meta-learning and few-shot adaptation, enabling robots to quickly adjust to new embodiments with minimal data, reducing the need for extensive retraining.
VI-D Real-Time Control
@@
VI-E Safety, Robustness, and Social Compliance
@@
Real-world dexterous manipulation presents significant challenges in error detection, recovery, and adaptability, requiring robots to operate reliably in dynamic and unstructured environments. Failure detection is complex due to sensor noise, occlusions, and unpredictable interactions, making it difficult to distinguish minor execution deviations from critical failures such as grasp failures or unexpected object motion. Once an error is detected, adaptive recovery strategies such as re-grasping and trajectory replanning must be computed in real time while maintaining stability and task continuity. Future research should address two key aspects. First, large-scale failure datasets and standardized benchmarks are essential for improving data-driven recovery policies. The lack of diverse, labeled failure cases across various objects, tasks, and environments limits model generalization. Establishing comprehensive datasets and evaluation protocols for failure detection, uncertainty estimation, and recovery effectiveness would provide a foundation for training and benchmarking robust policies. Second, self-supervised learning for multi-modal anomaly detection could enable robots to autonomously refine their error detection capabilities. By leveraging visual, tactile, and proprioceptive feedback, robots could learn to recognize and anticipate failures in real time, improving adaptability and robustness in dynamic environments.
@@
Safety is equally critical for both the robot and its surrounding environment, including human users, particularly in real-world deployments where unpredictable interactions and dynamic conditions pose significant risks. In dexterous manipulation, safety considerations involve collision avoidance, force regulation, and compliance control, particularly when interacting with fragile objects or operating near humans. However, achieving these safety measures requires handling varying contact conditions, but sensor noise, occlusions, and data processing delays can reduce reliability. Additionally, while compliant actuators and soft robotic designs help mitigate impact forces, integrating these hardware safety mechanisms involves trade-offs between control precision, responsiveness, and durability.
Beyond technical safety, social compliance is crucial for real-world deployment while less explored, particularly in human-robot interaction settings. Robots must adhere to social norms, ethical guidelines, and human expectations to be perceived as trustworthy and acceptable. This includes adapting manipulation strategies to align with human workspaces, ensuring transparent behavior and predictability, and minimizing actions that could cause discomfort or disruption. However, existing manipulation frameworks lack awareness of social constraints and human preferences. To address this challenge, interactive learning paradigms, where robots refine their socially compliant manipulation strategies by learning from human corrections and preferences, offer a promising research direction. Complementing this, multi-modal human-robot interaction datasets that integrate verbal, visual, and non-verbal cues would enhance contextual understanding, enabling robots to better anticipate and respond to human needs. Furthermore, ensuring consistency and reliability in socially aware dexterous manipulation requires standardized benchmarks for social compliance, providing objective evaluation criteria to assess how well robots integrate into human-centered environments.
@@ Acknowledgment
We would like to thank Qianyi Wang, Mingwu Liu, and Shouzheng Wang for their contributions to the completion of this survey.

### v4

@@ VI Challenges and Future Directions in Imitation Learning-Based Dexterous Manipulation
Imitation learning-based dexterous manipulation poses unique challenges due to the inherent complexities of both IL and dexterous control. Despite significant advancements over the past decade, several challenges hinder its human-level dexterity and real-world applicability. Figure 6 summarizes the challenges based on research impact and technical difficulty, highlighting immediate priorities that are further discussed in the remainder of this section.
@@ VI-A1 Heterogeneous Data Fusion
Dexterous manipulation relies on multi-modal sensory inputs (e.g., visual, tactile, proprioceptive, and force), each with varying sampling rates, noise characteristics, and spatial-temporal resolutions, making data integration and synchronization challenging. Moreover, differences in embodiments and gripper designs introduce additional complexities. For example, demonstrations collected with one robotic hand may not directly generalize well to another due to variations in kinematics, actuation mechanisms, and sensor placements. Addressing these challenges requires (1) multi-modal alignment techniques to improve sensor fusion and (2) cross-embodiment learning frameworks for better transferability across robotic platforms and embodiments.
@@ VI-A4 Data Collection Costs
Figure 6: Prioritization matrix of key challenges in imitation learning-based dexterous manipulation.
@@ VI-C Generalization to Novel Setups
Generalizing imitation learning-based dexterous manipulation policies is challenging due to task and environment variability, adaptive learning limitations, and cross-embodiment adaptability:
@@ VI-C2 Adaptive and Continual Learning Frameworks
VI-C3 Cross-Embodiment Adaptability
@@
Variability in robot embodiments, gripper designs, sensor configurations, and actuation dynamics poses significant challenges for generalization. A policy trained on one robotic hand may struggle to transfer to another due to differences in degrees of freedom, joint limits, contact dynamics, and control strategies. Even within the same robotic platform, inconsistencies arise from sensor noise, latency, and mechanical tolerances. To address this, morphology-agnostic policy learning can be explored, where models are trained across diverse robotic embodiments to develop transferable representations. Graph-based and latent-space embeddings of robot kinematics could help policies reason about different embodiments more effectively. Additionally, modular policy architectures, where separate components (e.g., perception, control, and adaptation modules) are fine-tuned independently, may enhance transferability. Another promising direction is meta-learning and few-shot adaptation, enabling robots to quickly adjust to new embodiments with minimal data, reducing the need for extensive retraining.
VI-D Sim-to-Real Transfer
Simulation provides a scalable and controlled environment for training dexterous manipulation policies. However, the sim-to-real gap remains a major obstacle to deploying IL models on real-world robotic systems. This gap arises from discrepancies in factors such as contact dynamics, sensor noise, actuation delays, and material properties, which are often inadequately modeled in simulation. The following subsection discusses key techniques and ongoing challenges related to sim-to-real transfer.
VI-D1 Domain Randomization
One widely adopted strategy for bridging the sim-to-real gap is domain randomization. Instead of replicating real-world conditions exactly, domain randomization introduces variability in simulation parameters. This exposure helps policies generalize to unseen real-world scenarios [276]. However, domain randomization requires manual tuning of parameter distributions, which demands expert knowledge to balance realism and variability. It also struggles to address unmodeled dynamics, such as non-linear material deformations or high-frequency contact interactions, which are crucial for dexterous manipulation. Therefore, future research could focus on combining domain randomization with complementary sim-to-real techniques, such as real-world fine-tuning or adversarial domain adaptation, to achieve more robust and reliable transfer.
VI-D2 Feature Alignment
Another promising line of work focuses on feature alignment, where the objective is to map simulated and real-world sensory inputs into a shared latent feature space. This can be achieved through representation learning techniques such as autoencoders or contrastive learning, which minimize the distributional discrepancy between the domains. Similarly, cross-modal embeddings can be learned to align tactile or proprioceptive features from both domains, which is particularly valuable for dexterous tasks involving high-dimensional state observations. These approaches can complement domain randomization by providing a structured way to reduce domain shift.
VI-D3 Adversarial Domain Adaptation
Adversarial techniques have also emerged as a powerful way to minimize sim-to-real discrepancies. This approach has been applied to tasks where direct collection of real-world demonstrations is expensive [327]. In robotic manipulation, adversarial adaptation can be used in combination with IL by first imitating expert demonstrations in simulation and then fine-tuning the learned policy to fool a domain discriminator, ensuring smoother deployment in real-world settings. Future research could explore integrating adversarial adaptation with generative models or self-supervised representation learning to improve robustness and scalability, particularly for high-dimensional sensory inputs such as vision and tactile feedback.
VI-D4 Hybrid Training and Online Adaptation
Beyond domain adaptation, hybrid training paradigms are gaining significant traction. In these approaches, policies are first pre-trained in simulation for rapid skill acquisition and subsequently fine-tuned on real-world data using IL, RL, or offline corrections [328]. This strategy combines the efficiency and scalability of simulation with the precision of real-world adaptation, thereby reducing the need for extensive real-world data collection. Additionally, self-supervised real-to-sim refinement, where real-world rollouts are used to iteratively adjust simulation parameters, can improve the fidelity of simulated environments and facilitate bi-directional transfer.
VI-E Real-Time Control
@@
VI-F Safety, Robustness, and Social Compliance
@@
Real-world dexterous manipulation demands reliable error detection, recovery, and adaptability in dynamic, unstructured environments. Detecting failures is challenging due to sensor noise, occlusions, and unpredictable interactions, while recovery strategies like re-grasping or trajectory replanning must be executed in real time to maintain stability and task continuity. Future research should address two key aspects. First, large-scale failure datasets and standardized benchmarks are essential for improving data-driven recovery policies. Establishing comprehensive datasets and evaluation protocols for failure detection, uncertainty estimation, and recovery effectiveness would provide a foundation for training and benchmarking robust policies. Second, self-supervised learning for multi-modal anomaly detection could enable robots to autonomously refine their error detection capabilities. By leveraging visual, tactile, and proprioceptive feedback, robots could learn to recognize and anticipate failures in real time, improving adaptability and robustness in dynamic environments.
@@
Safety is equally critical, particularly in real-world deployments where unpredictable interactions and dynamic conditions pose significant risks. In dexterous manipulation, safety considerations involve collision avoidance, force regulation, and compliance control, particularly when interacting with fragile objects or operating near humans. However, achieving these safety measures requires handling varying contact conditions, but sensor noise, occlusions, and data processing delays can reduce reliability. Additionally, while compliant actuators and soft robotic designs help mitigate impact forces, integrating these hardware safety mechanisms involves trade-offs between control precision, responsiveness, and durability.
To address safety constraints in IL for real-world dexterous manipulation, recent studies have investigated constrained policy optimization [329] and Lyapunov-based safe learning [330]. Constrained policy optimization integrates hard limits on torque, contact force, or joint velocity directly into the training objective, often via Lagrangian relaxation or primal-dual methods. For instance, Constrained Trust Region Policy Optimization (C-TRPO) [331] augments the trust-region update rule by incorporating actuator and contact-force limits, ensuring that each policy update remains close to expert demonstrations while strictly respecting physical safety bounds. In parallel, Lyapunov-based approaches synthesize control policies that provably satisfy safety specifications, constructing Lyapunov functions as certificates of stability. In dexterous manipulation, Neural Lyapunov Control (NLC) [332] has been adapted to learn neural network policies that maintain force closure and avoid excessive contact forces. Recent work combines these paradigms. Xiao et al. [333] propose SafeDiffuser, which augments diffusion-policy-based planning with control barrier functions to enforce real-time safety constraints. By conditioning the diffusion model on expert demonstrations and integrating CBF-based corrections at every denoising step, the method achieves robust generalization to unseen scenarios while guaranteeing safety-critical bounds.
Beyond technical safety, social compliance is essential for real-world deployment yet remains underexplored, particularly in HRI settings. In the context of dexterous manipulation, prior research has investigated various approaches, such as RL-based methods (e.g., preference-based RL[334] and inverse RL[335]), MPC frameworks (e.g., socially aware Model Predictive Path Integral control[336]), and game-theoretic models (e.g., Nash equilibrium formulations[337]). However, achieving social compliance remains challenging due to the inherent ambiguity of human comfort, spatial preferences, and social norms, which are often context-dependent and difficult to model explicitly. Additionally, soft social constraints, such as maintaining legibility of motion, avoiding intrusive actions, or signaling intent, are subjective and require learning from diverse human feedback, which is costly and time-consuming. Another challenge lies in the lack of standardized evaluation metrics and benchmarking protocols to assess social compliance. Existing metrics, such as trajectory smoothness, proximity compliance, or motion legibility, are often evaluated qualitatively or through small-scale user studies without a consistent framework. Moreover, there is a lack of publicly available datasets and simulation environments specifically designed for evaluating social compliance in dexterous manipulation, unlike traditional benchmarks for grasping or motion planning (e.g., RLBench [338]or ManiSkill3 [339]). Without well-defined metrics, datasets, and test scenarios, progress in this area remains fragmented and difficult to measure systematically.
VI-G High-Precision and Micro-Scale Manipulation
While the above discussion focuses on macro-scale object manipulation (e.g., bottles, tools), scaling down to the millimeter or micrometer level (e.g., electronic components, micro-machines) introduces fundamentally different challenges. At these scales, perception limitations, such as low sensor resolution, visual noise, and occlusion, become critical. Actuation requirements also rise sharply, demanding ultra-precise force and motion control within micrometer tolerances to prevent damage to delicate objects. Conventional actuators frequently suffer from backlash, hysteresis, and limited resolution, while achieving high control precision becomes a bottleneck. These challenges have driven the development of advanced control methods [340, 341, 342], including force-feedback, adaptive impedance, and hybrid position-force strategies designed specifically for micro-scale tasks.
Moreover, data collection at micro-scales remains a significant challenge. Acquiring large, diverse, and high-quality demonstrations is both costly and time-intensive, given the specialized equipment and delicate experimental setups required. Synthetic data generation through high-fidelity physics-based simulation is a promising direction, but it requires accurate modeling of micro-scale physical effects, such as adhesion forces, surface roughness, or fluid dynamics, which are often negligible in macro-scale tasks. From an algorithmic perspective, recent research [74] underscores the potential of hierarchical IL frameworks in addressing dexterous manipulation in micro-scale tasks. These techniques, together with advances in small-scale robots[343], present a promising pathway for enabling autonomous micro-manipulation tasks, such as precision assembly and surgical procedures.
@@ Acknowledgment
We would like to thank Huapeng Li, Qianyi Wang, Shouzheng Wang, and Mingwu Liu for their contributions to the completion of this survey.